### v4

@@ Abstract
….. In nonsmooth, nonconvex stochastic optimization, understanding the uniform convergence of subdifferential mappings is crucial for analyzing stationary points of sample average approximations of risk as they approach the population risk. Yet, characterizing this convergence remains a fundamental challenge. This work introduces a novel perspective by connecting the uniform convergence of subdifferential mappings to that of subgradient mappings as empirical risk converges to the population risk. We prove that, for stochastic weakly-convex objectives, and within any open set, a uniform bound on the convergence of subgradients—chosen arbitrarily from the corresponding subdifferential sets—translates to a uniform bound on the convergence of the subdifferential sets themselves, measured by the Hausdorff metric. Using this technique, we derive uniform convergence rates for subdifferential sets of stochastic convex-composite objectives. Our results do not rely on key distributional assumptions in the literature, such as the continuous differentiability of the population objective, yet still provide tight convergence rates. These guarantees lead to new insights into the nonsmooth landscapes of such objectives within finite samples.
@@ 1 Introduction
At the core of nonsmooth, nonconvex stochastic optimization lies a fundamental challenge—characterizing the statistical uniform convergence behavior of set-valued subdifferential mappings [1]. This characterization is not only key to understanding the convergence of stationary points defined by subdifferentials but also essential for analyzing statistical estimation and the performance of numerical algorithms. Consequently, understanding the uniform convergence of these subdifferential mappings for the empirical risk plays a pivotal role in stochastic programming [2, Section 7.2]. Nevertheless, existing results often rely on continuous differentiability assumptions for the population risk or fall short of providing tight convergence rates, underscoring the need to develop new analytic tools for characterizing the convergence behaviors of subdifferential sets (see Section 1.1 for detailed discussions on related works).
Towards addressing this gap, this work provides new understandings of the uniform convergence behavior of subdifferential sets by presenting a general principle that, notably, extends to settings without relying on the smoothness of population risks. Specifically, our work targets at stochastic weakly convex objectives—a prominent class of nonsmooth and nonconvex objectives with widespread applications [3, 4, 5]. Our approach hinges on a fundamental property of weakly convex functions. Specifically, for any pair of weakly convex functions, we establish that over any arbitrary open set within their domains, the uniform bound on the distance between their respective set-valued subdifferential mappings, measured via the Hausdorff metric, is upper bounded by the uniform bound on the distance between any corresponding vector-valued subgradient mappings of these functions (Theorem 1).  This deterministic result—which does not rely on differentiability—immediately implies a crucial connection between the uniform convergence of subdifferentials and the uniform convergence of subgradient mappings in sample average approximations, enabling the analysis of the former through the latter (Theorem 2).
@@
We further apply our technique to an important subclass of stochastic weakly-convex objectives called stochastic convex-composite objectives [6, 7], which include robust phase retrieval [8], blind deconvolution [9], matrix sensing [10, 11], conditional Value-at-Risk [12], and nonlinear quantile regressions [13], commonly encountered in the field of machine learning, statistics, imaging, and risk management (Theorem 3). Our analysis establishes uniform convergence of subdifferential sets for stochastic weakly convex objectives by studying the convergence behavior of stochastic subgradient mappings (Theorem 4 and Theorem 5). Our result, which does not rely on a key assumption in the literature requiring the population risk to be continuously differentiable, provides tight convergence rates in the context of many existing studies (see Section 1.1 for a detailed explanation). For example, in the setting of robust phase retrieval [8, 14], our results (Theorem 1) achieve a sharp rate of $\sqrt{d/m}$ (modulo logarithmic factors on $d,m$) for the uniform convergence of subdifferential mappings under the Hausdorff distance, where $d$ is the problem’s dimension and $m$ is the sample size. This is complementary to the state-of-the-art $4]{d/m}$ rate, which is established under the graphical distance of subdifferentials using an alternative approach based on Attouch’s epigraphical convergence theorem [15, 16]. Our result leads to new understanding of the nonsmooth landscape of these objectives, e.g., the landscape of the finite sample robust phase retrieval initially studied in [14], including the locations and convergence properties of stationary points.

### v5

@@ Abstract
….. We investigate the uniform convergence of subdifferential mappings from empirical risk to population risk in nonsmooth, nonconvex stochastic optimization. This question is key to understanding how empirical stationary points approximate population ones, yet characterizing this convergence remains a fundamental challenge due to the set-valued and nonsmooth nature of subdifferentials. This work establishes a general reduction principle: for weakly convex stochastic objectives, over any open subset of the domain, we show that a uniform bound on the convergence of selected subgradients—chosen arbitrarily from subdifferential sets—yields a corresponding uniform bound on the Hausdorff distance between the subdifferentials. This deterministic result reduces the study of set-valued subdifferential convergence to simpler vector-valued subgradient convergence. We apply this reduction to derive sharp uniform convergence rates for subdifferential mappings in stochastic convex-composite optimization, without relying on differentiability assumptions on the population risk. These guarantees clarify the landscape of nonsmooth empirical objectives and offer new insight into the geometry of optimization problems arising in robust statistics and related applications.
@@ 1 Introduction
At the core of nonsmooth, nonconvex stochastic optimization lies a fundamental challenge—characterizing the statistical uniform convergence behavior of set-valued subdifferential mappings [1]. This characterization is not only key to understanding the convergence of stationary points defined by subdifferentials but also essential for analyzing statistical estimation and the performance of numerical algorithms. Consequently, understanding the uniform convergence of these subdifferential mappings for the empirical risk plays a pivotal role in stochastic programming [2, Section 7.2]. Nevertheless, existing approaches to this problem often require the population risk to be continuously differentiable. Such smoothness assumptions can obscure the geometric structure of nonsmooth models and limit the applicability of theoretical guarantees in real-world problems (see Section 1.1 for detailed discussions on related works).
In this work, we provide a new perspective on this problem by establishing a general deterministic principle that links the uniform convergence of subgradient selections to the uniform convergence of subdifferential sets. Specifically, we show that for any pair of weakly convex functions, the uniform distance between their subdifferential mappings (measured in Hausdorff distance) is bounded above by the uniform distance between any pair of subgradient selections over any open domain (Theorem 1). This result bypasses any smoothness assumptions and allows us to characterize subdifferential set convergence by analyzing simpler vector-valued subgradient mappings. Importantly, it enables the analysis of the uniform convergence of subdifferentials in stochastic optimization settings via the uniform convergence of subgradients, leading to sharp statistical rates.
@@
We instantiate this principle in the context of stochastic weakly convex objectives, a broad class of nonsmooth, nonconvex functions encompassing many widely studied models [3, 4, 5]. Our main application is to stochastic convex-composite objectives [6, 7], which include robust phase retrieval [8], blind deconvolution [9], matrix sensing [10, 11], conditional Value-at-Risk [12], and nonlinear quantile regressions [13]. These models appear across imaging, statistics, machine learning, and risk management. For this class, we derive uniform convergence rates for subdifferentials by studying the convergence behavior of stochastic subgradient mappings (Theorem 4 and Theorem 5). Our analysis complements prior work by (i) removing differentiability assumptions on the population risk, and (ii) establishing tight convergence rates under the Hausdorff distance, a topologically weaker notion than graphical convergence that more directly captures the geometry of subdifferential sets. For instance, in the setting of robust phase retrieval [8, 14], we establish a sharp convergence rate of $\sqrt{d/m}$ (modulo logarithmic factors on $d,m$) for the uniform convergence of subdifferential mappings under the Hausdorff distance, where $d$ is the problem’s dimension and $m$ is the sample size (Corollary 1). This is complementary to the state-of-the-art $4]{d/m}$ rate, which is established under the graphical distance of subdifferentials using an alternative approach based on Attouch’s epigraphical convergence theorem [15, 16]. Our results lead to new understanding into nonsmooth empirical risk landscapes, including how the locations of stationary points evolve with sample size—providing a sharper understanding of finite-sample behavior in high-dimensional stochastic optimization.